QUESTION 11 Find the solution of x + 2x +x=f(t), x(0)=1, x(o=0, where f(t) = 1 if t< 2; and f(t) = 0 if t> 2.
Answer #1

A solution given x +Qx! + x = Elt), X10 )= 1.X110: 0 Whenne fit)=(1 if t22 ♡ o if t 22 0 01 => Ft) 2 dx + X + dt² dt equatioD+1=0 & D+ 120 - D = -1:-1 MJs-dif Mas -1 300ts of quxilhoy equn @ Os some Corresponding part of cofo Pe of the form as Cie mSo P. I. 1 eot 024qxott P.IE 1 . eot 0 +0+1 P.IS A {eote 14 When Fic)= 0. if tzo then P. I oo 1 D2420+1 P. I = 0 B Hence subsFrom eguh ③ we get X; t Oxe + t Gei{ +1 Cgt e-t at X il n diff. equh. 0 wor. to. t, we get X Co [ ette-t] to Cg [ e-tatet usFrom equh we get X 1. e-t t X e-t + Gg te-te a te-t - © . 7 diff equn © w. r. to t we gel xs Let + ( [et-tet? →x=net + C2O= -1 teq Cq z 1 X = e-t t dit ect 07, if flt=0.622 X e-te t.te-t Hence complete Solution using Boundlahy Condition, we eget.